A population has a mean of 200 and a standard deviation of 60. Suppose a sample of size 125 is selected and x-bar is used to estimate µ . Use z-table.

a.) What is the probability that the sample mean will be within ±3 of the population mean (to 4 decimals)?

b.) What is the probability that the sample mean will be within ±14 of the population mean (to 4 decimals)?
Answer #1

Here we have mu = 200, sigma = 60, n = 125

a)  The probability that the sample mean will be within ±3 of the population mean :

p ( -3< overline{x}-mu<3 )

S.E = sigma /sqrt{n} = 60/V125 = 5.37

So ,

p ( -3< overline{x}-mu<3 )=pleft ( rac{-3}{5.37}<rac{overline{x}-mu }{sigma /sqrt{n}}<rac{3}{5.37} ight )

= p ( -0.56 < z < 0.56 )

= p ( z < 0.56 ) - p ( z < -0.56 )

= 0.7123 - 0.2877

= 0.4246

b)

The probability that the sample mean will be within ±14 of the population mean :

p ( -14< overline{x}-mu<14 )

S.E = sigma /sqrt{n} = 60/V125 = 5.37

So ,

p ( -14< overline{x}-mu<14 )=pleft ( rac{-14}{5.37}<rac{overline{x}-mu }{sigma /sqrt{n}}<rac{14}{5.37} ight )

= p ( -2.61 < z < 2.61 )

= p ( z < 2.61 ) - p ( z < -2.61 )

= 0.9955 - 0.0045

= 0.9910